# PICore: Physics-Informed Unsupervised Coreset Selection for Data Efficient Neural Operator Training

### v1

@@ Abstract
Neural operators offer a powerful paradigm for solving partial differential equations (PDEs) that cannot be solved analytically by learning mappings between function spaces. However, there are two main bottlenecks in training neural operators: they require a significant amount of training data to learn these mappings, and this data needs to be labeled, which can only be accessed via expensive simulations with numerical solvers. To alleviate both of these issues simultaneously, we propose PICore, an unsupervised coreset selection framework that identifies the most informative training samples without requiring access to ground-truth PDE solutions. PICore leverages a physics-informed loss to select unlabeled inputs by their potential contribution to operator learning. After selecting a compact subset of inputs, only those samples are simulated using numerical solvers to generate labels, reducing annotation costs. We then train the neural operator on the reduced labeled dataset, significantly decreasing training time as well. Across four diverse PDE benchmarks and multiple coreset selection strategies, PICore achieves up to $78\%$ average increase in training efficiency relative to supervised coreset selection methods with minimal changes in accuracy. We provide code at https://github.com/Asatheesh6561/PICore.
@@ 1 Introduction
Operator learning has emerged as a tool for accelerating PDE solutions by developing data-driven approximations using neural networks instead of traditional grid-based discretizations. Neural operators (Kovachki et al., 2023) are a family of neural networks that learn mappings between function spaces, such as initial conditions to solutions, which allows for resolution-invariant predictions. Models such as Fourier Neural Operator (FNO) (Li et al., 2020) and U-Net Neural Operator (UNO) (Rahman et al., 2023) have shown state-of-the-art performance on various PDE benchmarks, and the ability to generalize to higher-order resolutions with minimal performance drops. Additional work, such as Physics Informed Neural Operator (PINO) (Li et al., 2024c) and Markov Neural Operator (MNO) (Li et al., 2021), incorporates additional losses into neural operator training to improve performance and increase convergence speed.
@@
Coreset selection (Agarwal et al., 2005; Sener & Savarese, 2017) is a data-efficient training strategy that identifies a subset of the original training data that is most informative for model learning. Once this subset is identified, training only needs to be done on this subset, significantly reducing training time. However, this requires the full labeled training data to select a subset, which does not alleviate the cost of collecting labels. On the other hand, active learning (Gu et al., 2021; Cao & Tsang, 2022) minimizes data annotation costs by only labeling a subset of the training data at each iteration. Active learning selects a subset by a proxy metric such as Bayesian (Zhao et al., 2021; Beluch et al., 2018) or representation-based methods (Yang & Loog, 2022; Kim & Shin, 2022) at each training iteration, and trains only on that subset. While this does decrease labeling costs at each iteration, active learning requires a significant portion of the dataset for training, reducing convergence speed (Li et al., 2024a). Thus, we pose the following research question:
@@
We address this problem using unsupervised coreset selection by identifying the most informative training samples based on the physics-informed loss (Li et al., 2024c), a criterion that does not require any ground truth labels. By leveraging this loss, we can prioritize samples likely to improve model performance without the need for expensive simulations. Ground truth labels are then generated only for this selected subset, significantly reducing the overall annotation cost. Finally, we train neural operator models on the reduced, high-quality dataset, leading to faster training times without compromising accuracy.
@@
  • We present the first comprehensive benchmark for coreset selection in the context of neural operator learning. Through extensive experiments on four diverse PDE datasets, we show that PICore achieves competitive accuracy to supervised methods while dramatically improving end-to-end training efficiency by up to 78% relative to supervised coreset selection and 5×5$\times$5 × relative to non-coreset baselines.
@@ 2.1 Neural Operators
2.2 Coreset Selection
@@
The traditional testbed for coreset selection algorithms has been image classification tasks, but it also has applications in Neural Architecture Search (NAS) (Shim et al., 2021), efficient GAN training (Sinha et al., 2020), continual learning (Yoon et al., 2021), and large language model (LLM) finetuning (Zhang et al., 2025). However, to the best of our knowledge, coreset selection has not been used for improving the training efficiency of neural operator learning.
@@ 2.3 Data Efficiency for Neural Operators
The closest existing work to our own is Chen et al. (2024), which develops an unsupervised pretraining strategy that leverages Masked Autoencoders (MAEs) to learn effective unsupervised representations, which are then used to fine-tune with a smaller ground-truth dataset. However, this indirectly addresses issues with training efficiency and data labeling costs using a two-stage training process, whereas PICore directly addresses both problems in a single training cycle. Hemmasian & Farimani (2024) avoid running expensive simulations on high-resolution data by pretraining neural operators in low dimensions, but this requires a factorized neural operator such as Factorized Fourier Neural Operator (FFNO) (Tran et al., 2021). In contrast, our method is independent of the operator architecture. Li et al. (2024b) uses an active learning strategy to reduce labeling costs from running simulations by maximizing a utility cost ratio. However, this is specific to FNO and only addresses the cost of data annotation and not training efficiency.
@@ 4 PICore
17:     $u_{i}^{\dagger}\leftarrow\mathcal{G}^{\dagger}(a_{i})$triangleright Run numerical simulation
18:     $D_{c}\leftarrow D_{c}\cup\{(a_{i},u_{i}^{\dagger})\}$
@@
Before coreset selection, we warm-start the neural operator with the physics-informed loss over the full dataset for a small number of epochs $T_{w}<<T$ Warm starting is common in prior coreset selection methods (Killamsetty et al., 2021a) and is necessary as most coreset selection algorithms require gradient information, which is unusable with a randomly initialized model. We provide the full algorithm in Algorithm 1.
@@ 5 Experimental Details
Additional information on the datasets can be found in Section A. We use the Fourier Neural Operator (FNO) (Li et al., 2020) and U-Net Neural Operator (Rahman et al., 2023) as the base models for all experiments due to their implementation simplicity and performance. However, PICore can work out of the box with any neural operator. We also use 5 coreset selection algorithms in our experiments: CRAIG (Mirzasoleiman et al., 2020), GradMatch (Killamsetty et al., 2021a), AdaCore (Pooladzandi et al., 2022), EL2N (Paul et al., 2021) and graNd (Paul et al., 2021). Additional information on these coreset selection algorithms can be found in Section B. We use coreset selection percentages of 20%, 30%, 40%, 60%, and 80%.
@@
We use $T_{w}=25$ warmup epochs and reset the neural operator to its initialization to ensure fair comparisons between supervised and physics-informed coreset selection. Then, we train neural operators for $T=500$ epochs and report the average NRMSE over 5 seeds on a held-out test set at the input resolution. All experiments were run on a single RTX A4000 GPU.
@@ 6.1 Main Results
We report the core findings for PICore and supervised coreset selection across the four representative PDE datasets in Tables 1, 2, 3, and 4. In addition to the average test NRMSE over all coreset selection algorithms, we show the decrease in full training time (including data annotation costs through simulation) relative to the non-coreset selection baseline. Our results demonstrate that PICore consistently achieves competitive test performance compared to supervised coreset selection while providing substantial computational efficiency gains, primarily by reducing expensive data annotation (simulation) costs during the coreset selection phase.
PICore significantly improves training efficiency through reduced simulation costs.
Across four representative PDE datasets—Advection, Burgers, Darcy, and Navier-Stokes Incompressible—PICore consistently reduces the total training time by cutting down expensive simulation-based annotation. These efficiency gains become especially significant as the complexity of the PDE increases: Across the four datasets, PICore achieves average training time reductions of 0.9%, 9.8%, 30.1%, and 78.0% compared to supervised coreset selection, calculated by averaging the relative acceleration improvements at each selection percentage (20%, 30%, 40%, 60%, and 80%). For example, at a 20% coreset size, PICore achieves a 5.01$\times$ speedup on Darcy Flow (vs. 2.24$\times$ for supervised methods) and a 5.00$\times$ speedup on Navier-Stokes (vs. 1.14$\times$) using UNO. These results show that PICore scales well to high-dimensional scientific problems where data annotation costs dominate training.
@@
Despite the substantial efficiency improvements, PICore remains competitive with strong supervised baselines in terms of test NRMSE. For instance, at a 20% coreset size, PICore achieves NRMSE values of $3.77\times 10^{-2}$ for Advection using FNO and $3.42\times 10^{-2}$ for Burgers using UNO—very close to the supervised method values of $3.69\times 10^{-2}$ and $3.32\times 10^{-2}$, respectively. This holds across different coreset sizes, with minor variations, showing that PICore can be as effective as supervised coreset selection, even at low coreset selection percentages. However, we observe that not all coreset selection algorithms perform equally as well. Due to the convexity assumptions and Hessian approximations with CRAIG (Mirzasoleiman et al., 2020) and AdaCore (Pooladzandi et al., 2022), they have higher NRMSE losses compared to the other algorithms.
There is a tradeoff between efficiency and absolute test accuracy.

### v2

@@ Abstract
Neural operators offer a powerful paradigm for solving partial differential equations (PDEs) that cannot be solved analytically by learning mappings between function spaces. However, there are two main bottlenecks in training neural operators: they require a significant amount of training data to learn these mappings, and this data needs to be labeled, which can only be accessed via expensive simulations with numerical solvers. To alleviate both of these issues simultaneously, we propose PICore, an unsupervised coreset selection framework that identifies the most informative training samples without requiring access to ground-truth PDE solutions. PICore leverages a physics-informed loss to select unlabeled inputs by their potential contribution to operator learning. After selecting a compact subset of inputs, only those samples are simulated using numerical solvers to generate labels, reducing annotation costs. We then train the neural operator on the reduced labeled dataset, significantly decreasing training time as well. Across four diverse PDE benchmarks and multiple coreset selection strategies, PICore achieves up to $78\%$ average increase in training relative to supervised coreset selection methods with minimal changes in accuracy.
@@ 1 Introduction
Operator learning has emerged as a tool for accelerating PDE solutions by developing data-driven approximations using neural networks instead of traditional grid-based discretizations. Neural operators (Kovachki et al., 2023) are a family of neural networks that learn mappings between function spaces, such as initial conditions to solutions, which allows for resolution-invariant predictions. Models such as Fourier Neural Operator (FNO) (Li et al., 2020) and U-Net Neural Operator (UNO) (Rahman et al., 2023) have shown state-of-the-art performance on various PDE benchmarks, and the ability to generalize to higher-order resolutions with minimal performance drops. Additional work, such as Physics Informed Neural Operator (PINO) (Li et al., 2024c) and Markov Neural Operator (MNO) (Li et al., 2021b), incorporates additional losses into neural operator training to improve performance and increase convergence speed.
@@
Coreset selection (Agarwal et al., 2005; Sener & Savarese, 2017) is a data-efficient training strategy that identifies a subset of the original training data that is most informative for model learning. Once this subset is identified, training only needs to be done on this subset, significantly reducing training time. However, this requires the full labeled training data to select a subset, which does not alleviate the cost of collecting labels. On the other hand, active learning (Gu et al., 2021; Cao & Tsang, 2022) minimizes data annotation costs by only labeling a subset of the training data at each iteration. Active learning selects a subset by a proxy metric such as Bayesian (Zhao et al., 2021; Beluch et al., 2018) or representation-based methods (Yang & Loog, 2022; Kim & Shin, 2022) at each training iteration, and trains only on that subset. A limitation of many iterative active learning strategies is that repeatedly alternating between selecting points and updating the model can increase training time and reduce convergence speed. Thus, we pose the following research question:
@@
We address this problem using unsupervised coreset selection by identifying the most informative training samples based on the physics-informed loss (Li et al., 2024c), a criterion that does not require any ground truth labels. Our approach can also be viewed as a single-shot active learning implementation, where a subset of points is selected in one pass rather than iteratively. By leveraging this loss, we can prioritize samples likely to improve model performance without the need for expensive simulations. Ground truth labels are then generated only for this selected subset, significantly reducing the overall annotation cost. Finally, we train neural operator models on the reduced, high-quality dataset, leading to faster training times without compromising accuracy.
@@
  • We present the first comprehensive benchmark for coreset selection in the context of neural operator learning. Through extensive experiments on four diverse PDE datasets, we show that PICore achieves competitive accuracy to supervised methods while dramatically improving end-to-end training by up to 78% relative to supervised coreset selection.
@@ 2.1 Neural Operators
2.2 Data Efficient Machine Learning
2.2.1 Coreset Selection
@@
The traditional testbed for coreset selection algorithms has been image classification tasks, but it also has applications in Neural Architecture Search (NAS) (Shim et al., 2021), efficient GAN training (Sinha et al., 2020), continual learning (Yoon et al., 2021), and large language model (LLM) finetuning (Zhang et al., 2025). However, to the best of our knowledge, coreset selection has not been used for improving the training of neural operator learning.
2.2.2 Active Learning
In contrast to coreset selection, active learning, over multiple iterations and in an unsupervised environment, chooses previously unannotated data to label and trains on those newly labeled pairs. The key differences are that active learning is unsupervised, choosing training samples with only features and that active learning is done over many iterations instead of in a single shot. Many algorithms transfer from coreset selection to active learning. Aside from the equivalent random selection, there are cluster based methods for active learning to find representative and typical examples (Sener & Savarese, 2017; Hacohen et al., 2022) and uncertainty based methods that find data for which the model is either uncertain or degraded (Rahmati et al., 2024; Houlsby et al., 2011; Ash et al., 2019). Musekamp et al. (2024) shows several uncertainty based sampling methods and active learning methods for physics informed learning, but these are limited to PINNs and not to neural operators, which are more powerful due to their ability to map between function spaces, but are inherently more difficult to perform active learning.
@@ 2.3 Data Efficiency for Neural Operators
The closest existing work to our own is Chen et al. (2024), which develops an unsupervised pretraining strategy that leverages Masked Autoencoders (MAEs) to learn effective unsupervised representations, which are then used to fine-tune with a smaller ground-truth dataset. However, this indirectly addresses issues with training and data labeling costs using a two-stage training process, whereas PICore directly addresses both problems in a single training cycle. Hemmasian & Farimani (2024) avoid running expensive simulations on high-resolution data by pretraining neural operators in low dimensions, but this requires a factorized neural operator such as Factorized Fourier Neural Operator (FFNO) (Tran et al., 2021). In contrast, our method is independent of the operator architecture. Li et al. (2024b) uses an active learning strategy to reduce labeling costs from running simulations by maximizing a utility cost ratio. However, this is specific to FNO and only addresses the cost of data annotation and not training efficiency. Finally, (Kim et al., 2025) uses a surrogate model for active learning, but this only increases efficiency at the timestep level while ours can reduce the data annotation cost of full solutions in a single shot.
@@ 4 PICore
[…]
@@
Before coreset selection, we warm-start the neural operator with the physics-informed loss over the full dataset for a small number of epochs $T_{w}<<T$. Warm starting is common in prior coreset selection methods (Killamsetty et al., 2021a) and is necessary as most coreset selection algorithms require gradient information, which is unusable with a randomly initialized model. We provide the full algorithm in Algorithm 1.
@@ 5 Experimental Details
Each dataset has 1000 generated trajectories, with 900 that can be used for training (varying based on the coreset selection percentage) and 100 for testing, which is comparable to existing neural operator literature (Li et al., 2021a; 2024c). We generate 20 timesteps forward for Advection and Burgers, and only 10 timesteps for the Navier Stokes Incompressible dataset due to memory limits. Additional information on the datasets can be found in Section A. We use the Fourier Neural Operator (FNO) (Li et al., 2020) and U-Net Neural Operator (Rahman et al., 2023) as the base models for all experiments due to their implementation simplicity and performance. However, PICore can work out of the box with any neural operator. We also use 5 coreset selection algorithms in our experiments: CRAIG (Mirzasoleiman et al., 2020), GradMatch (Killamsetty et al., 2021a), AdaCore (Pooladzandi et al., 2022), EL2N (Paul et al., 2021) and graNd (Paul et al., 2021). CRAIG, AdaCore, and GradMatch are submodular methods that try to match the gradient sum of the coreset to the gradient sum of the entire dataset. GraNd and EL2N are score based methods that use the gradient or the loss. Additional information on these coreset selection algorithms can be found in Section B. We use coreset selection percentages of 20%, 30%, 40%, 60%, and 80%.
@@
We use $T_{w}=25$ warmup epochs and reset the neural operator to its initialization to ensure fair comparisons between supervised and physics-informed coreset selection. Then, we train neural operators for $T=500$ epochs and report the average NRMSE over 5 seeds on a held-out test set at the input resolution. We calculate the acceleration as the total time taken for supervised coreset selection / PICore (including data generation, warm starting, and training time) divided by the total time for the non-coreset baseline.
In addition to supervised coreset selection, we compare PICore to random subset selection, pure unsupervised training with the physics informed loss, and an active learning baseline based on uncertainty. Since most active learning baselines are for classification problems, we extend loss-as-uncertainty methods in Liu & Li (2023); Beluch et al. (2018) to neural operators. For the active learning baseline, we begin by randomly selecting 10% of the available data as an initial training set and generating the corresponding ground-truth PDE solutions. We then train 10 independent copies of the neural operator on this subset for $T_{w}$ epochs, which is the same training time for the other methods. After training, we construct the final coreset from the remaining unlabeled data in a single step by selecting the points exhibiting the highest variance across the model predictions.
@@ 6.1 Main Results
We report the core findings for PICore and supervised coreset selection across the four representative PDE datasets in Tables 1, 2, 3, and 4. In addition to the average test NRMSE over the best coreset selection algorithm for each method, we show the decrease in full training time (including data annotation costs through simulation) relative to the non-coreset selection baseline. Our results demonstrate that PICore consistently achieves competitive test performance compared to supervised coreset selection while providing substantial computational gains, primarily by reducing expensive data annotation (simulation) costs during the coreset selection phase.
PICore significantly improves training through reduced simulation costs.
[…]
As shown in Tables 6, 6, 8, and 8, the relative contributions of training and data generation speedups vary by dataset difficulty. For simpler datasets such as Advection and Burgers, gains are driven primarily by reductions in training time. For example, Advection achieves a 79.7% improvement in training time but only a 1.47% improvement in data generation time at the 20% coreset level. In contrast, for more challenging datasets, the impact of training time reductions diminishes, while reductions in data generation time play a more significant role in overall gains. These results show that PICore scales well to high-dimensional scientific problems where data annotation costs dominate training.
@@
Despite the substantial improvements, PICore remains competitive with strong supervised baselines in terms of test NRMSE. For instance, at a 20% coreset size, PICore achieves NRMSE values of $3.46\times 10^{-2}$ for Advection using FNO and $2.84\times 10^{-2}$ for Burgers using UNO, very close to the supervised method values of $3.42\times 10^{-2}$ and $2.93\times 10^{-2}$, respectively. This holds across different coreset sizes, with minor variations, showing that PICore can be as effective as supervised coreset selection, even at low coreset selection percentages. In fact, many dataset, model, and selection percentages combinations show that PICore improves upon supervised coreset selection, but there is largely no significant change in accuracy between the two methods. However, we observe that not all coreset selection algorithms perform equally as well. Due to the convexity assumptions and Hessian approximations with CRAIG (Mirzasoleiman et al., 2020) and AdaCore (Pooladzandi et al., 2022), they have higher NRMSE losses compared to the other algorithms. Thus, the tables almost always report GradMatch, GraNd, or El2N as the coreset algorithm to use for PICore and supervised coreset selection.
Coreset Selection methods outperform Random and Active Learning baselines on most datasets.
Random subset selection consistently underperforms relative to both PICore and subset selection, and this difference increases as we increase the complexity of the dataset and decrease the selection percentage. For example, random selection has an nRMSE of $2.74\times 10^{-1}$ on the Navier Stokes Incompressible dataset at a 20% coreset selection percentage, where as PICore has an nRMSE of $1.12\times 10^{-2}$ and Supervised Coreset Selection has an nRMSE of $9.57\times 10^{-2}$. This shows that using PDE specific information (either supervised loss or the physics informed loss) is necessary to achieve a more accurate solution with less data. Interestingly, we see that active learning outperforms the PICore on the Advection dataset with UNO on medium coreset selection percentages (40-60%) and on the Navier Stokes Incompressible dataset with FNO. However, it is much worse on all other dataset and model combinations by a considerable margin.
Physics-informed training alone is insufficient to achieve high accuracy.
Training a neural operator solely with the physics-informed loss on the full unlabeled dataset yields significantly worse performance than all other baselines, despite requiring no simulation cost. For example, on the Navier–Stokes Incompressible dataset with FNO, the physics-informed-only baseline achieves an nRMSE of $1.46\times 10^{-0}$, compared to PICore’s $1.25\times 10^{-1}$ at 20% coreset selection, a decrease of an order of magnitude. This large gap arises because the physics-informed loss is unstable as a standalone training objective and fails to capture fine-grained solution details without supervised guidance. In contrast, PICore leverages the physics-informed loss as a proxy for selecting informative samples, then trains on a small labeled subset with the supervised nRMSE loss, which mitachieving both higher accuracy and stability.
There is a tradeoff between and absolute test accuracy.